given a population of 2000 normally distributed stores with mean equal to 74, and a standard deviation of 6. How many scores are between 50 and 70?


1
Expert's answer
2022-05-05T11:01:12-0400
"n=2000, \mu=74,\sigma=6,Z=\frac{(x-μ)}{σ}""P(50<X<70)=P(Z<\frac{70-74}{6})"




"-P(Z\le\frac{50-74}{6})"




"=0.25249-0.00003"




"=0.25246"



"0.25246(2000)=505"


"505" scores